OpenStudy (anonymous):

Which function below represents the arithmetic sequence 3, 7, 11, 15...? f(n) = 4 + 3(n – 1) f(n) = 4 + 3n f(n) = 3 + 4n f(n) = 3 + 4(n – 1)

OpenStudy (anonymous):

i can help on this one if you'd like

OpenStudy (anonymous):

Sure, thanks

OpenStudy (anonymous):

you can see that its just adding 4 to every number correct?

OpenStudy (anonymous):

pretty much yeah

OpenStudy (anonymous):

so with every number 4 will be added, hence the variable n

OpenStudy (anonymous):

yea..

OpenStudy (anonymous):

still can't figure it out?

OpenStudy (anonymous):

i can i can i can i can i can i can i can i an i can i can now i can./...

OpenStudy (anonymous):

ok what is it ?

OpenStudy (anonymous):

i think its D

OpenStudy (anonymous):

hint: the variable is next to 4

OpenStudy (anonymous):

mmmmm try again

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

oops sorry i got it mixed up d is right

OpenStudy (anonymous):

@severus4ever You are correct. It is the last choice.

OpenStudy (anonymous):

WHOOHOO

OpenStudy (anonymous):

i saw a negative sign somewhere

OpenStudy (anonymous):

Watch...I will show you the method...so every time you have an arithmetic sequence, you will be able to write the function on your own (even without choices)....

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

For an arithmetic sequence, use a+ (n-1)d where a = first term, d = common difference In this sequence 3,7,11,15,.... d = common difference = 4 a= first term = 3 so we have: 3 + (n-1)(4) which is the last choice. It can be simplified to 3 + 4n if you distribute and combine like terms.

OpenStudy (anonymous):

thank you so much!

OpenStudy (anonymous):

Sorry... 3+ (n-1)(4) = 3 + 4n - 4 = -1 + 4n OR 4n -1

OpenStudy (anonymous):

thanks again, sorry to bother again..but can you also help me with a little science?

OpenStudy (anonymous):

@severus4ever Welcome. I only in mathematics.

OpenStudy (anonymous):

I am only in mathematics.

OpenStudy (anonymous):

ok thanks anyways